OpenStudy (anonymous):

what is the one of the real number values that must be exluded from the domain of the function

OpenStudy (anonymous):

\[g(x)=\frac{ 3x-6 }{ x ^{2}-7x +12 } ?\]

jimthompson5910 (jim_thompson5910):

hint: solve for x x^2 - 7x + 12 = 0

OpenStudy (anonymous):

x=4?

jimthompson5910 (jim_thompson5910):

that's one solution, what's another

OpenStudy (anonymous):

but i only need one.... and i dont know others.... i substitute x from 1-8 :3 and got 4

jimthompson5910 (jim_thompson5910):

smart thinking, plug in values to see which ones give you zero another way is to factor and use the zero product property to solve

OpenStudy (anonymous):

whats the zero product property?

jimthompson5910 (jim_thompson5910):

if A*B = 0, then A = 0 or B = 0

jimthompson5910 (jim_thompson5910):

so if you factored x^2 - 7x + 12 = 0 to get (x-3)(x-4) = 0

jimthompson5910 (jim_thompson5910):

you can use the zero product property to go from (x-3)(x-4) = 0 to x-3 = 0 or x - 4 = 0

jimthompson5910 (jim_thompson5910):

then you solve each equation

OpenStudy (anonymous):

so x= 4 or x=3

jimthompson5910 (jim_thompson5910):

yep, those two values make the denominator equal to 0, so they must be excluded

OpenStudy (anonymous):

oh